# Semantic Latent Motion for Portrait Video Generation

### v1

@@ Abstract
Recent advancements in portrait video generation have been noteworthy. However, existing methods rely heavily on human priors and pre-trained generation models, which may introduce unrealistic motion and lead to inefficient inference. To address these challenges, we propose Semantic Latent Motion (SeMo), a compact and expressive motion representation. Leveraging this representation, our approach achieve both high-quality visual results and efficient inference. SeMo follows an effective three-step framework: Abstraction, Reasoning, and Generation. First, in the Abstraction step, we use a carefully designed Mask Motion Encoder to compress the subject’s motion state into a compact and abstract latent motion (1D token). Second, in the Reasoning step, long-term modeling and efficient reasoning are performed in this latent space to generate motion sequences. Finally, in the Generation step, the motion dynamics serve as conditional information to guide the generation model in synthesizing realistic transitions from reference frames to target frames. Thanks to the compact and descriptive nature of Semantic Latent Motion, our method enables real-time video generation with highly realistic motion. User studies demonstrate that our approach surpasses state-of-the-art models with an 81% win rate in realism. Extensive experiments further highlight its strong compression capability, reconstruction quality, and generative potential. Moreover, its fully self-supervised nature suggests promising applications in broader video generation tasks.
@@ 1 Introduction
Figure 1: Semantic Latent Motion is a compact and expressive representation of the character’s current state. The motion dynamics in the semantic latent space serve as a powerful guiding condition, steering the generative model from the reference image to the target image. Leveraging the compactness of latent motion, we can model long sequences and perform more efficient resoning.
Figure 2: Overview of SeMo. The proposed framework follows a three-step process (abstraction, reasoning, generation) implemented in two training stages. In the first stage, we train a Motion AutoEncoder for abstraction and generation. During the abstraction step, image features are subjected to random masking and compression to obtain a compact and abstract latent motion. In the generation step, latent motion is used as a condition to guide the diffusion decoder in generating the target image from the reference image. In the second stage, we train a Motion Generator for reasoning, which generates motion sequences from audio.
@@
Existing approaches mainly follow two paths. The first path integrates human prior knowledge of motion representation, such as sequences of 3DMM, pose, landmarks, etc. as expressions of motion [9, 4], combined with a generative renderer to produce the final video frames. The second path utilizes pre-trained video generation models, using their strong video generation capabilities and conditional control of cross-attention to learn the mapping from audio to video frames [6, 4, 41]. Many methods combine both of these paths. However, in the first path, human prior knowledge relies on domain-specific heuristics, often failing to capture detailed motions, which results in human videos that lack naturalness and coherence. Moreover, this framework is not adaptable to multiple scenarios (for example, it cannot be used for general video generation, as it is impractical to extract pose sequences for all objects in the scene). In the second path, the complex spatiotemporal attention modeling and multi-step denoising in diffusion models prevent real-time generation and long-context modeling.
@@
We observe that these motion representations ( 3DMM, pose, landmarks ) all express a high-level semantic:“ the positions and attributes of meaningful parts.” They are all abstractions of human knowledge about facial dynamics and are limited by design biases. For instance, we place landmarks on faces because we believe that changes in the positions of the noise, eyes, and mouth corners are crucial for facial dynamics. Similarly, we use skeletons because we believe that skeletal structures are important for body movements. However, these representations are limited by human abstracting capabilities and cannot capture detailed motion. We ask: Can deep learning models abstract a semantic latent representation that better captures detailed face motions by self-supervised learning? We will introduce our motivation from the following two aspects.
1) Representing video motion in a compact and expressive space. Modeling video motion presents significant challenges. Fortunately, facial videos exhibit more constrained motion ranges and patterns compared to general videos. Previous works have demonstrated that models can synthesize realistic images by extracting detailed information from reference images and semantic information from prior knowledge. Human-crafted priors (e.g., facial landmarks) represent compact and semantic motion representations abstracted through domain knowledge. If humans can achieve this, deep learning models could potentially learn superior representations through self-supervision – capturing more compact and detail-preserving motion characteristics. This motivates our exploration of learning semantic latent motion representations that surpass human-designed priors in expressiveness.
Figure 3: Given an audio clip and a reference image, SeMo can generate lifelike video. From top to bottom, we highlight several key features of our model: a. The generated characters exhibit natural head movements and blinking while speaking. b. They display rich motion dynamics with substantial amplitude. c. Model can adapt to different poses. d. Model generalize well across various datasets. e. The generated videos feature remarkably natural lighting and shadow variations.
@@
2) Employing self-supervised semantic representations to guide generation. Current generative models [5, 17, 1, 2, 43, 23, 46, 13] predominantly rely on human-abstracted conditions like text prompts or landmarks. While effective, these representations are inherently limited by human abstraction capabilities and domain-specific biases, often failing to capture subtle motion details. Recent works like RCG [18] and StoryDiffusion[52] have demonstrated that self-supervised latent features can serve as superior generation conditions compared to manual priors. Building on this insight, we propose using learned semantic latent motion as conditional signals for video generation, enabling more precise and detailed motion control.
Motivated by these insights, we design a universal motion representation framework coupled with a simple yet effective video generation pipeline. Our approach consists of two key components: First, we train a Motion Autoencoder, where an encoder compresses each facial video frame into a compact latent motion representation (1D token) using a random masking strategy. The diffusion-based decoder then reconstructs the target frames using reference images as inputs and semantic latent motion as conditional signals. Once the Motion Autoencoder is trained, we extract latent motion representations for downstream tasks. For portrait video generation, we train a diffusion-based generator that synthesizes latent motion sequences directly from audio inputs. The compact latent space enables efficient long-term temporal modeling and preserves comprehensive event information. During inference, we first generate motion sequences from audio and reference images, then synthesize video frames using the Motion Autoencoder’s decoder.
The rigorous reconstruction task encourages our latent motion representation to develop strong descriptive power, effectively capturing essential information within an extremely compact space. Leveraging this representation, our method achieves highly natural and vivid animations, outperforming state-of-the-art methods with an FID score of 26.8 (compared to the 27.4 baseline) and achieving an 81% win rate in realism in a user study.
Beyond these results, we further observe that even when 90% of the video frame content (input to the motion encoder) is masked, our Motion AutoEncoder can still generate coherent and smooth videos. This reveals two critical insights: 1) High-level semantic information sufficiently guides video generation, 2) Video motion modeling requires less information proportion than image generation (contrasting 75% masking ratio in MAE [14]). Moreover, as a general framework, our Motion AutoEncoder demonstrates strong adaptability to generic videos containing limited subjects and simple motions, such as object rotations.
@@ 2 Related Works
Motion-driven video synthesis techniques typically generate target frames by combining reference frames with motion representations. FOMM [30] established a baseline through sparse keypoint detection and affine transformation estimation to derive dense motion fields. LIA [39] advanced this paradigm by decomposing latent vectors and employing optical flow-based warping operations. MRAA [31] extended the framework by separately predicting foreground region motions and background displacements, subsequently rendering through pixel-wise flow manipulation. IMF [10] introduced a fully self-supervised approach that encodes reference and target frames into compact latent tokens. This method progressively reconstructs target frame pixels through a coarse-to-fine process that integrates reference image features at multiple granularity levels. These developments collectively demonstrate the evolution from explicit motion parameterization towards more implicit, learned representations that better preserve spatial-temporal coherence.
@@
The portrait generation task has garnered significant attention in recent years due to its broad applications and has witnessed rapid advancements. Wave2Lip [25] pioneered the use of GAN networks combined with lower-face masking to directly synthesize video frames from reference images and audio inputs. Subsequent works including VividTalk [33], SadTalker [48], and Moda [22] have leveraged 3D morphable priors coupled with neural rendering techniques to generate target video frames. Notably, VASA-1 [42] and Gaia [15] introduced hybrid architectures that encode both 3D prior knowledge and appearance features into a latent space, establishing a more universal and compact motion representation framework. Next, diffusion-based generators are employed to produce motion sequences conditioned on audio inputs [34, 19, 32, 49]. Recent innovations like EMO [34] and Hallo2 [6] have further enhanced visual quality through the integration of pre-trained video generation foundation models with specialized reference networks. AniTalker [21] presents a two-stage self-supervised learning paradigm. The overarching trend in this domain demonstrates a clear shift from hand-crafted priors towards data-driven approaches that enable models to capture finer details and enhance the naturalness of synthesized visuals.
@@
A large amount of self-supervised learning work  [14, 28, 35, 11, 26, 44, 37] has utilized encoded features for various downstream visual understanding tasks. Self-supervised representations have been proven to possess rich semantics and can serve as powerful guidance signals. In RCG  [18], the quality of image generation was significantly improved by introducing features obtained through MoCo’s [3] self-supervised learning framework. Meanwhile, Story Diffusion [52] obtains a continuous guidance signal by performing linear interpolation in the latent space of CLIP features, thereby controlling the video generation model to produce more coherent content.
@@ 3 Method
Figure 4: Illustration of Motion AutoEncoder
Figure 5: An example of video reconstruction: In the example above, we demonstrate that SeMo can not only accurately model facial movements but also capture environmental changes. With a token size of 1, semantic latent motion is not only more compact than the motion in other works but also captures more comprehensive information.
Given a reference image and its corresponding audio, our goal is to generate a dynamic video with synchronized motion and sound. To achieve this, we propose a two-stage self-supervised learning framework. In the first stage, the objective is to extract a compact, high-level semantic motion $\mathbf{M}\in\mathbb{R}^{l\times d}$ for each frame of the image. We first use learnable tokens and random masking to compress the motion into an abstract representation, and then utilize the reference image and the motion to reconstruct the target image. Once the Motion AutoEncoder is trained, it is frozen. In the second stage, we train a Motion Generator. For each frame, we first use the motion encoder to extract the motion and then generate the target motion based on the audio and the corresponding motion of the reference image. The details of these two stages will be described below.
@@ 3.1 Motion AutoEncoder
Motion Encoder. Our motion encoder design is motivated by the observation that facial movements - akin to 3D Morphable Models (3DMM) and keypoint trajectories - can be effectively represented in a highly compact space. We propose a novel Masked Motion Encoder architecture that integrates dual strategies of random masking and latent space compression to derive semantically rich latent motion.
@@
Specifically, we first compress input frames $\mathbf{I}\in\mathbb{R}^{H\times W\times C}$ into latent representations $\mathbf{X}\in\mathbb{R}^{h\times w\times c}$ using a pre-trained stable diffusion VAE encoder [29]. Following ViT [8], we embed each latent feature into non-overlapping patches $\mathbf{Z}\in\mathbb{R}^{t\times d}$, where $t=\frac{H}{p}\times\frac{W}{p}$ denotes the token count of each image, $p$ denotes the patch size, and apply random masking with ratios uniformly sampled from 0% to 90%. The masking ratios are carefully designed, low mask ratios are used to encourage the model to retain more complete information, while high mask ratios aim to enhance the model’s ability to abstract. Once the masking is complete, we discard the masked tokens and retain the unmasked tokens, arranging them into a sequence in their original order. As shown in Fig. 10, we demonstrate that even with 90% of the content masked, the model is still able to generate acceptable results.
$\small\mathbf{M}=\operatorname{Enc}(\mathbf{Z^{m}}\oplus\mathbf{T})$ (1)
where $\oplus$ denotes concatenation, $\operatorname{Enc}$ denotes Motion Encoder, $\mathbf{Z^{m}}$ denotes masked image patches.
@@
TikTok [45] has demonstrated that image features can be represented as 1D sequence representations. Therefore, we designed a 1D learnable token sequence, which is adjustable in length and can become very compact. As shown in Fig. 4, we concatenate the learnable tokens with the image features and process them through ViT layers. Ultimately, latent motion $\mathbf{M}$ is output at the position corresponding to the learnable tokens. Here, the number of motion tokens and the token dimension size are consistent with the learnable tokens. Note that, unless otherwise specified, our latent motion is set to 1 token with a dimension of 512. Tab. 5 Tab. 6 varies number and dimension of tokens.
Diffusion Decoder. At the stage, our objective is to use the reference image as a reference for detailed information, with motion serving as the semantic condition, guiding a generative model to fill in missing details. Our decoder is a standard DiT [24] model. Specifically, we randomly select two sets of images from the data, obtaining reference image $\mathbf{Z_{r}}$ and target image feature $\mathbf{Z_{t}}$. We apply the same masking ratio to image features and obtain reference motion $\mathbf{M_{r}}$ and target motion $\mathbf{M_{t}}$ through the motion encoder. Following RectifiedFlow[20], We add noise to the unmasked target image feature Eq. 2, and concatenate noised target image feature $\mathbf{Z_{t}^{n}}$ with the $\mathbf{Z_{r}}$, $\mathbf{M_{r}}$ and $\mathbf{M_{t}}$ Eq. 3. Finally, the DiT decoder outputs the predicted noise velocity $\hat{\mathbf{V}}$. We refer to the module described above as the Transfer Block.
$\small\mathbf{Z_{t}^{n}}=(1-t)\mathbf{Z_{t}}+t\epsilon$ (2)
$\small\hat{\mathbf{V}}=\operatorname{Dec}(\mathbf{Z_{r}}\oplus\mathbf{Z_{t}^{n}}\oplus\mathbf{M_{r}}\oplus\mathbf{M_{t}})$ (3)
Where $\epsilon$ denotes Gaussian Noise sampled from standard normal distribution, $t\in[0,1]$ denotes the intensity of the noise, $\mathbf{Z_{t}^{n}}$ denotes noisy target image feature.
The above module enables the transfer from a reference image to a target image. We can extend this framework to transfer from the reference image to target video. Specifically, we randomly select one reference frame and a sequence of $f$ consecutive video frames. After applying Transfer Block to each frame, we introduce a Temporal Align Block. This is because we observed that treating each frame independently for denoising results in a lack of smooth transitions in the details. In Temporal Align Block, We flatten noised target image feature $\mathbf{Z^{n}_{t}}\in\mathbb{R}^{f\times t\times d}$ to $\mathbf{Z^{n}_{t}}\in\mathbb{R}^{b\times d}$, where $f$ denotes number of frames, $b=f\times t$. After transformation, we apply self-attention for temporal modeling, The complexity of this block is $o(f)$, which does not lead to significant computational overhead.

### v2

# A Self-supervised Motion Representation for Portrait Video Generation

@@ Abstract
Recent advancements in portrait video generation have been noteworthy. However, existing methods rely heavily on human priors and pre-trained generative models, Motion representations based on human priors may introduce unrealistic motion, while methods relying on pre-trained generative models often suffer from inefficient inference. To address these challenges, we propose Semantic Latent Motion (SeMo), a compact and expressive motion representation. Leveraging this representation, our approach achieve both high-quality visual results and efficient inference. SeMo follows an effective three-step framework: Abstraction, Reasoning, and Generation. First, in the Abstraction step, we use a carefully designed Masked Motion Encoder, which leverages a self-supervised learning paradigm to compress the subject’s motion state into a compact and abstract latent motion (1D token). Second, in the Reasoning step, we efficiently generate motion sequences based on the driving audio signal. Finally, in the Generation step, the motion dynamics serve as conditional information to guide the motion decoder in synthesizing realistic transitions from reference frame to target video. Thanks to the compact and expressive nature of Semantic Latent Motion, our method achieves efficient motion representation and high-quality video generation. User studies demonstrate that our approach surpasses state-of-the-art models with an 81% win rate in realism. Extensive experiments further highlight its strong compression capability, reconstruction quality, and generative potential.
@@ 1 Introduction
Figure 1: Comparison of Motion Representation and Generation Methods. Motion representation: (a) Explicit motion representation (e.g., 3DMM, facial landmarks). (b) Implicit motion representation guided by explicit priors (e.g., implicit keypoints). (c) Our self-supervised implicit representation compresses motion into a compact 1D token without pretrained models or prior knowledge. Video generation approaches: (d) Adapter-modulated generation fine-tunes a pretrained video model via driving-signal adapters. (e) Identity-motion feature wrapping fuses identity features with target motion. (f) Our method leverages reference and target motion as conditions for efficient generation.
@@
Existing approaches mainly follow two paths. In the first path (fig. 1e), methods directly utilize explicit motion representations, or employ explicit priors to learn an implicit motion representation; these motion representations are then fused with identity features to generate video frames. In the second path (fig. 1d), pre-trained video generation models are used, utilizing their powerful generative capacity and conditional cross-attention mechanisms to learn the mapping from the driving signal to video frames [6, 4, 42]. Many methods combine both of these paths. However, in the first path, human prior knowledge relies on domain-specific heuristics, often failing to capture detailed motions, which results in human videos that lack naturalness and coherence. Moreover, this framework is not adaptable to multiple scenarios (e.g., 3DMM is only applicable to facial scenarios and cannot be generalized to other types of scene). In the second path, the computationally intensive spatial-temporal attention operations - with their quadratic scaling in memory and computation - combined with multistep denoising in diffusion models create significant bottlenecks for video generation efficiency.
@@
We observe that these motion representations (3DMM, pose, landmarks) all express a high-level semantic:“ the positions and attributes of meaningful parts.” They are all abstractions of human knowledge about facial dynamics and are limited by design biases. For instance, we place landmarks on faces because we believe that changes in the positions of the noise, eyes, and mouth corners are crucial for facial dynamics. Similarly, we use skeletons because we believe that skeletal structures are important for body movements. However, these representations are limited by human abstracting capabilities and cannot capture detailed motion. We investigate the fundamental question: How can we derive a motion representation that effectively captures fine-grained facial dynamics without relying on prior knowledge? To address this, we first present our motivation from two key perspectives.
1) Video motion in a highly compact and semantic space. Prior works [34, 49, 40] have demonstrated that the content between frames in general videos contains substantial redundancy, video motion can be compressed into a very small latent space. For portrait videos in particular, the motion can be represented even more compactly, as the motion range of human subjects is inherently limited and the background typically remains static. Previous approaches leverage human-crafted semantic priors (such as 3DMM or landmarks) to represent motion through domain knowledge, and have validated the effectiveness of this strategy. This suggests that, by surpassing the abstraction capacity and compression limitations of human-designed priors, it may be possible to encode portrait video motion into an even more compact and semantic latent space.
@@
2) Self-supervised representations as guidance. Current generative models [5, 16, 1, 2, 44, 21, 50, 12] predominantly rely on human-abstracted conditions like text prompts or landmarks. While effective, these representations are inherently limited by human abstraction capabilities and domain-specific biases, often failing to capture subtle motion details. Recent works like RCG [17] and StoryDiffusion[56] have demonstrated that self-supervised latent features can serve as superior generation conditions compared to manual priors. Building on this insight, we propose using learned semantic latent motion as conditional signals for video generation.
Motivated by these insights, we propose a universal motion representation framework paired with a simple yet effective video generation pipeline. Our approach comprises two key stages: In the first stage, we train a Motion Autoencoder in which the encoder compresses each video frame into a compact latent motion representation (1D token) using a random masking strategy. This encourages the model to abstract relevant semantic motion while discarding irrelevant details. The rectified flow-based decoder then leverages its powerful generative capabilities to fill in visual details, conditioned on the reference image and reference motion. In the second stage, we train a standard DiT model to directly generate motion sequences from audio and reference motions. Thanks to the highly compact nature of our motion representation, this process is extremely efficient. Our motion extraction and video generation pipeline is both simple and straightforward to generalize to other scenarios, requiring neither pretrained models nor hand-crafted human priors. Benefiting from the strong descriptive power of our semantic latent motion representation, our method produces highly natural and vivid animations, surpassing state-of-the-art approaches with an FID score of 26.8 (compared to a baseline of 27.4) and achieving an 81% win rate in realism according to user studies. In summary, our contributions to the community include:
  1. 1.
    We present a new representation framework that fuses the semantic descriptive capability of self-supervised representations with the fine-grained generative fidelity of a rectified-flow model, offering a new perspective on dynamic representation.
  2. 2.
    We show that the proposed representation encodes rich, complex motion cues in an compact latent space, leading to better reconstruction quality in video reconstruction tasks.
  3. 3.
    We validate the effectiveness of our approach in audio-driven portrait video generation. Owing to the self-supervised nature of SeMo, which eliminates the need for pretrained models or manually-designed priors, our method adapts well to a wide range of challenging scenarios while maintaining strong performance.
@@ 2 Related Works
Motion-driven video synthesis techniques typically generate target frames by combining reference frames with motion representations. FOMM [29] established a baseline through sparse keypoint detection and affine transformation estimation to derive dense motion fields. LIA [39] advanced this paradigm by decomposing latent vectors and employing optical flow-based warping operations. MRAA [30] extended the framework by separately predicting foreground region motions and background displacements, subsequently rendering through pixel-wise flow manipulation. IMF [9] introduced a fully self-supervised approach that encodes reference and target frames into compact latent tokens. The work most closely related to ours is IMF [9]appendix E illustrates the key differences between our approach and IMF.
@@
The portrait generation task has garnered significant attention in recent years due to its broad applications and has witnessed rapid advancements. Wave2Lip [23] pioneered the use of GAN networks combined with lower-face masking to directly synthesize video frames from reference images and audio inputs. Subsequent works including VividTalk [32], SadTalker [52], and Moda [20] have leveraged 3D morphable priors coupled with neural rendering techniques to generate target video frames. Notably, VASA-1 [43] and Gaia [14] introduced hybrid architectures that encode both 3D prior knowledge and appearance features into a latent space, establishing a more universal and compact motion representation framework. Next, generative models are employed to produce motion sequences conditioned on audio inputs [33, 18, 31, 53]. Recent innovations like EMO [33] and Hallo2 [6] have further enhanced visual quality through the integration of pre-trained video generation foundation models with specialized reference networks. AniTalker [19] presents a two-stage self-supervised learning paradigm.
@@
A large amount of self-supervised learning work  [13, 26, 35, 10, 24, 45, 37] has utilized encoded features for various downstream visual understanding tasks. Self-supervised representations have been proven to possess rich semantics and can serve as powerful guidance signals. In RCG  [17], the quality of image generation was significantly improved by introducing features obtained through MoCo’s [3] self-supervised learning framework. Meanwhile, Story Diffusion [56] obtains a continuous guidance signal by performing linear interpolation in the latent space of CLIP features, thereby controlling the video generation model to produce more coherent content. REPA [48] and VA-VAE [46] demonstrate the ability to improve generation and reconstruction by aligning latent features with strong self-supervised visual representations.
Figure 2: Overview of SeMo. The proposed framework follows a three-step process (abstraction, reasoning, generation) implemented in two training stages. In the first stage, we train a Motion AutoEncoder for abstraction and generation. During the abstraction step, image features are subjected to random masking and compression to obtain a compact and abstract latent motion. In the generation step, latent motion is used as a condition to guide the motion decoder in generating the target video from the reference image. In the second stage, we train a Motion Generator for reasoning, which generates motion sequences from audio and reference motions.
@@ 3 Method
Given a reference image and its corresponding audio, our goal is to generate a dynamic video with synchronized motion and sound. To achieve this, we propose a two-stage self-supervised learning framework. In the first stage, the objective is to extract a compact, high-level semantic motion $m\in\mathbb{R}^{l\times d}$ for each frame. We first use learnable tokens and random masking strategy to compress the motion into an abstract representation, and then utilize the reference image, reference motion and target motion to reconstruct the videos. In the second stage, we train a Motion Generator and the Motion AutoEncoder is frozen. For each frame, we first use the motion encoder to extract the motion and then generate motion sequence based on the audio and reference motion. The details of these two stages will be described below.
@@ 3.1 Motion AutoEncoder
Our core motivation is to leverage self-supervised learning to extract the most salient motion information, and then utilize the powerful distribution modeling ability of rectified flow to progressively complete the high-frequency details conditioned on the motion information in a coarse-to-fine manner.
@@
Motion Encoder. Our motion encoder design is motivated by the observation that facial movements - akin to 3D Morphable Models (3DMM) and keypoint trajectories - can be effectively represented in a highly compact space. We propose a new Masked Motion Encoder architecture that integrates dual strategies of random masking and latent space compression to derive semantically rich latent motion. Specifically, we first compress input frames $I\in\mathbb{R}^{H\times W\times C}$ into latent representations $X\in\mathbb{R}^{h\times w\times c}$ using a pre-trained stable diffusion VAE encoder [27]. Following ViT [8], we embed each latent into non-overlapping patches, and apply random masking with ratios uniformly sampled from 0% to 90%. Once the masking is complete, we discard the masked tokens and retain the unmasked tokens, arranging them into a sequence in their original order. The masking ratios are carefully designed: low mask ratios encourage the model to retain more complete information, while high mask ratios are intended to strengthen the model’s ability to ignore fine-grained details and focus on preserving the most salient motion information. As shown in fig. 6, applying a masking strategy can greatly enhance the model’s generalization ability by ignoring unnecessary information. The mask strategy can enhance the model’s generalization ability fig. 10 and also accelerate convergence during the second-stage training fig. 7. More analysis is provided in fig. 7.
@@
TikTok [47] has demonstrated that image features can be represented as 1D sequence representations. Therefore, we designed a 1D learnable token $y\in\mathbb{R}^{l\times d}$ where $l$ denotes the number of tokens and $d$ denotes the features dimension. we concatenate the learnable token with the image features and process them through ViT layers. Ultimately, latent motion $M$ is output at the position corresponding to the learnable tokens. Here, the number of motion tokens and the token dimension size are consistent with the learnable tokens. Each image is encoded independently; we refer to the image features and motion associated with the reference image as $x_{\text{ref}}$,$m_{\text{ref}}$, and those associated with the target image as $x_{\text{tar}}$,$m_{\text{tar}}$. The encoding process can be formulated as follows: $m_{\text{ref}}=\operatorname{Enc}_{\theta}(x_{\text{ref}})$ , $m_{\text{tar}}=\operatorname{Enc}_{\theta}(x_{\text{tar}})$, where $\theta$ is the parameter of the motion encoder.
Figure 3: Motion Decoder. The Motion Control Module uses the reference motion $m_{\text{ref}}$ and target motion $m_{\text{tar}}$ as control signals. The goal of this module is to transfer from the reference image features $x_{\text{ref}}$ to the target image features $x_{\text{tar}}$. After reshaping, the target image features $x_{\text{tar}}$ are processed with self-attention along the temporal dimension, enabling alignment of high-frequency information.
Motion Decoder. At the decoding stage, our objective is to use the reference image for appearance information, with motion serving as the semantic condition, guiding a generative model to fill in missing details. Our decoder consists of two modules: the Motion Control Module and the Temporal Align Module (MCM and TAM for short). In the Motion Control Module, we concatenate the reference image features $x_{\text{ref}}$, the noised target image features $\bar{x}_{\text{tar}}$, the reference motion $m_{\text{ref}}$, and the target motion $m_{\text{tar}}$, then process them through a self-attention layer followed by a feed-forward network (FFN). The noising process follows the schedule in rectified flow, which can be formulated as: $\bar{x}=(1-t)x+t\epsilon$, where $t\in[0,1]$ is the noise level, and $\epsilon$ denotes Gaussian Noise sampled from standard normal distribution. $z_{\text{tar}}=\operatorname{MCM}_{\psi_{1}}(x_{\text{ref}},\bar{x}_{\text{tar}},m_{\text{ref}},m_{\text{tar}})$. Since denoising process in Motion Control Module is performed on each target frame, there could be inconsistencies in the high-frequency details across target frames. Therefore, in the Temporal Align Module, we concatenate the target image features $\{z_{\text{tar}}^{f}\}_{f=1}^{k}$ along the temporal dimension and apply self-attention to align their high-frequency information, where $k$ is the number of frames. $\{x_{\text{tar}}^{f}\}_{f=1}^{k}=\operatorname{TAM}_{\psi_{2}}(\{z_{\text{tar}}^{f}\}_{f=1}^{k})$. In all, the decoding process can be formulated as follows: $\{x_{\text{tar}}^{f}\}_{f=1}^{k}=\operatorname{Dec}_{\psi}(x_{\text{ref}},\{\bar{x}_{\text{tar}}^{f}\}_{f=1}^{k},m_{\text{ref}},\{m_{\text{tar}}^{f}\}_{f=1}^{k})$, where $\psi$ is the parameter of the motion decoder.